OpenStudy (ayee_ciera):

Which is an equation of a circle with center (2, -1) that passes through the point (3, 4)? (x + 2)2 + (y - 1)2 = 13 (x - 2)2 + (y + 1)2 = 26 (x - 2)2 + (y + 1)2 = 13 (x + 2)2 + (y - 1)2 = 26

OpenStudy (alexandervonhumboldt2):

(x-a)^2+(y-b)^2=r^2 where (a, b) center r radius

OpenStudy (alexandervonhumboldt2):

distance from center to (3, 4) is sqrt26.

OpenStudy (alexandervonhumboldt2):

so radius is sqrt26

OpenStudy (alexandervonhumboldt2):

can you write it now?

OpenStudy (ayee_ciera):

i dont get it

OpenStudy (alexandervonhumboldt2):

ok (x-a)^2+(y-b)^2=r^2 where (a, b) center r radius substitute (x-2)^2+(y+1)^2=(sqrt26)^2

OpenStudy (alexandervonhumboldt2):

(x - 2)2 + (y + 1)2 = 26

OpenStudy (alexandervonhumboldt2):

medal plz?